# A Generative Framework for Probabilistic, Spatiotemporally Coherent Downscaling of Climate Simulation

### v3

@@ Discussion
We introduced a probabilistic approach to joint spatial and temporal downscaling of multiple variables from climate to weather scale. The presented framework allows for using a trained generative model, which acts as a physical prior, in combination with an observation model that is formulated independent of the model training, and which defines the inference task. For our purpose, we leveraged this mechanism to encode a direct relationship between climate simulations and generated weather dynamics, such that the predicted small-scale phenomena preserve the statistics of the climate simulations. Our results provide a methodological advancement into joint spatial and temporal downscaling of multiple variables from ESM simulations to the weather scale.
High spatial and temporal resolution climate information is essential for impact assessment and decision-making on local scales, particularly in sectors sensitive to climate variability [12]. For instance, our approach can generate spatially and temporally consistent, multivariate scenarios that provide insights for critical infrastructure sectors such as agriculture, water resources, and renewable energy [13]. We briefly sketch out an example of the many potential downstream tasks by demonstrating the value of local climate information to estimate wind energy potentials [see for example 10]. We find that, locally, observational wind speed distributions differ substantially from the coarse ESM data, which can result in large differences in predicted wind power generation and accumulated yearly wind energy output. In the supplementary material (figure A.7), we illustrate that the downscaled predictions preserve the globally-averaged generated wind power, while locally, the estimates move closer to the reanalysis data. Thereby, the model corrects the predicted wind power when under- and over-estimated ESM wind power occurs. Beyond that, estimating small-scale temporal information, which preserves long-term global properties of the guiding ESM simulation, could prove useful for studying different types of extreme events, for example, heat waves [14, 15, 16, 17] or flooding events and hazards [18, 19, 20].
Current approaches to providing small-scale climate simulations concentrate on downscaling spatial snapshots for single variables [for example 21, 22, 23, 24, 25] and thus output de-coupled sequences of temporally independent states. Our approach closes this gap in that the jointly sampled multivariate predictions are each spatially, temporally, and physically coherent weather-scale climate trajectories. Further, existing work focuses on mapping between different reanalysis data sets [26, 27]. Our approach differs because we explicitly developed a model for unpaired samples that maps between different climate scenarios from ensembles from the CMIP6 project to realistic local weather patterns.
The generative model used in this study separates a trained physical prior model from the observations model; the latter specifies the inference task, depending on the application at hand. Thus, unlike approaches that require training the model directly on the conditioning information, the presented framework allows the formulation of explicit—and varying—functional relationships between observations and predictions. We see high future potential in encoding additional, potentially more complex, problem-specific domain knowledge through the observation model in subsequent studies that build upon the present work. For instance, one could encode known functional relationships between the predicted climate variables and external forcing, such as CO2 concentrations. It is a feature of the modular composition of the framework (figure 1) that the trained score model can be directly re-used in follow-up research. Trained weights, the model, and the implementations of the methods and experiments are provided at https://github.com/schmidtjonathan/Climate2Weather.
@@
Sampling spatiotemporally consistent trajectories requires increased computational cost compared to alternative approaches. While we have extended the original method by Rozet and Louppe [6] in terms of scalability, enabling it to process substantially longer trajectories, the temporal and spatial extent of the study is still limited by computational demands. This study limits itself to a small region with a diverse range of orography to establish and validate the methodological framework. However, we believe that long-range teleconnections that affect the sub-region under study are, to some extent, reflected in the predictions, given that the training data has been drawn from a Europe-wide reanalysis dataset. This argument is visually supported by embedding high-resolution predictions into a larger spatial context of reanalysis data in the supplementary figures A.3 to A.6.
The presented framework provides a flexible and powerful foundation for inference tasks in meteorology and climate science by separating between (1) capturing patterns from the learned dynamics and (2) imposing constraints through flexible and cheap conditioning. The model thus provides a step towards making long-term coarse-grained climate projections available for studying the impacts of anthropogenic climate change on local and multiple temporal scales.
@@ Methods
We analyze four closely interrelated atmospheric variables, namely the zonal (uas) and meridional (vas) components of near-surface wind speeds estimated at $10$ meters height, surface air temperature (tas), and sea level pressure (psl) due to their crucial importance for understanding atmospheric dynamics. They evolve on different spatial and temporal scales, enabling insights into the performance of downscaling methods across different scales.
Reanalysis data
For training (2006-2013) and evaluation (2014) of the model, we use data from the Consortium for Small-Scale Modeling Reanalysis 6 (COSMO-REA6) product that is a high-resolution reanalysis product for the European domain developed by the Deutscher Wetterdienst (DWD) [28] with a spatial resolution of approximately $6\leavevmode\nobreak$km and hourly temporal resolution. It serves as the ground truth observational dataset in the on-model evaluation of the presented methodological framework. The COSMO-REA6 data contains errors for some variables and lacks some observations for the years prior to 2006. In this work, only data from 2006 onwards was used.
Climate model data
We apply our downscaling model to historical model runs from two established general circulation models (GCMs) that are part of the sixth phase of the Coupled Model Intercomparison Project (CMIP6) [2], namely the higher-resolution earth system model (ESM) of the Max–Planck Institute (MPI-ESM1.2-HR) [29] and the high-resolution configuration of the third Hadley Centre Global Environment Model in the global coupled configuration 3.1 (HadGEM3-GC3.1-LM) [30]. Both models have an approximately $100\leavevmode\nobreak$km spatial and 6-hourly temporal resolution.
We restrict our analysis to the spatial subregion including parts of Germany, Switzerland, Austria and Czech Republic ($6^{\circ}\,\mathrm{E}-16^{\circ}\,\mathrm{E}$, $46^{\circ}\,\mathrm{N}-52^{\circ}\,\mathrm{N}$ (see supplementary figure A.1) which results in $128\times 128$ gridpoints per time point for the reanalysis dataset. We use data for the period 2006-2014, the time range in which the reanalysis and the historical GCM runs overlap. The GCM data is spatially re-gridded to a rotated latitude-longitude grid using bilinear interpolation to match the coordinates and grid type of the reanalysis dataset.
@@
The core of the presented methodological framework is a score-based diffusion model (DM) [31, 3, 4, 32, 33], an instance of generative models that involves training a deep neural network on a finite set of data points and can then be used to generate samples from the underlying data distribution. The basic formulation of DMs involves two main components. Firstly, a forward diffusion process transports the data distribution to a known, tractable distribution, such as, a standard normal distribution [3, 4]. In this process, structured data points (like RGB images, weather variables, etc.) are successively perturbed until all signals are entirely replaced by noise. The pivotal insight is that a certain class of diffusion processes are known to have a time-reversed counterpart, providing a generative process that successively refines samples from the noise distribution into structured data. The reversal of the diffusion process requires access to the score function, which can be understood as a function that, for any given degree of perturbation during the reverse-time diffusion process, separates noise from the signal to remove it. In the following, we will give a brief overview of the general framework and the extensions to it necessary to obtain the results presented in this work.
@@
The state $X(\tau)$ at $\tau=0$ is a data point $X_{0}\sim\mathcal{D}$ selected from a data set $\mathcal{D}$. The datum is successively perturbed through the diffusion process and thus loses all structure as $\tau\rightarrow\mathrm{T}$. Note that $\tau$ is sometimes referred to as "time", which does not mean physical time but rather a continuous degree of perturbation of the initial state. The drift $\mathrm{F}$ and dispersion $\mathrm{L}$ define functional properties of the forward process, which is driven by Brownian motion $W(\tau)$. At a final time step $\mathrm{T}$ (often $\mathrm{T}=1$), the process converges to a Gaussian distribution such that $X(\mathrm{T})\sim\mathcal{N}\left(0,\mathrm{\Pi}\right)$, where the final-step covariance $\mathrm{\Pi}$ depends on the choice of $\mathrm{F}$ and $\mathrm{L}$ and is often modeled to be the identity matrix. From a result by Anderson [35] the reverse process of equation 1 is known and given as
@@
This process is driven by reverse-time Brownian motion $\overleftarrow{W}(\tau)$ and depends on the gradient of the log-marginal density, called the score, of the diffused state $X(\tau)$ at every perturbation-time point $\tau$. We call equation 2 the generative process as it allows sampling unseen data points from the data distribution simply by
@@
    numerically simulating equation 2 backwards through time, starting from $X(\mathrm{T})$ and ending at a generated sample $X(0)$.
There is much existing and active research regarding the sampling algorithm used in step 2., which, in general, is slower and more complex when comparing diffusion models to other methods from the generative-model class [36, 37, 38]. This work uses a standard technique that solves an ordinary differential equation related to the marginal distribution of equation 2, instead [5, 4, 33].
In practice, the score of $p_{\tau}(X(\tau))$ is not accessible and has to be estimated. It is common practice to train a time-dependent neural network $s_{\theta}(X(\tau),\tau)$ on a finite set of samples from the data distribution to approximate the true, intractable score function. Most practical approaches optimize a de-noising score-matching objective [39, 3, 5, 4, 32, 33]. In light of this, simulating equation 2 is commonly perceived as iteratively de-noising the initial Gaussian random state $X(\mathrm{T})$ and successively refining it to a noise-free data point. For spatial data, a time-conditioned U-Net [40, 3] architecture has proven effective in modelling the score function. The U-Net architecture consists of multiple levels, each containing blocks of convolutional layers and skip connections. Along these levels, the spatial dimensionality of the input data is first successively reduced and then increased again. This encoder-decoder structure results in a bottleneck layer that forces the model to extract a limited set of meaningful features from the input. At each level, skip connections between equal-resolution blocks are introduced to facilitate the learning task by maintaining information throughout the encoding- and decoding process. This work uses a U-Net architecture with a total number of around $72$ million parameters, including a self-attention mechanism at the $8\times 8$-bottleneck layer [3, 41].
The framework used in the presented experiments is centered around the work by Rozet and Louppe [6] that extends the basic DM framework, enabling robust sampling from a posterior distribution and generating arbitrary-length sequences of spatiotemporally coherent state trajectories. The idea behind score-based data assimilation (SDA) is to separate an inference task into two parts:
@@
    learning a generative physical model that reproduces spatiotemporal patterns from the desired output space and
@@
    conditioning the prior model using a functional constraint (the observation model) to solve a specific task.
@@
In our case, the output space is that of high-resolution, multivariate spatiotemporal weather dynamics, so we train the physical model on reanalysis data. Regarding I., the score model is trained on and learns to generate short temporally and spatially coherent trajectories. Through an intricate convolution-like routine, the windows are assembled into trajectories of arbitrary length. De-noising each time step in the trajectory is computed using a short temporal context window around that time step. This yields output dynamics that are temporally coherent without the need for training the network on long sequences, which would not only significantly increase the computational complexity but also fix the number of time steps as a pre-determined hyperparameter, greatly limiting flexibility. We refer to Rozet and Louppe [6, Algorithm 2] for details and pseudo-code.
@@
For our experiments, we require scaling the original SDA framework to allow for the prediction of very long time horizons with thousands of steps. Unfortunately, sampling a spatiotemporally coherent trajectory in the described manner comes at an increased computational cost and, in particular, significant memory demands since the entire state trajectory must be held in memory. To make the approach scale to predicting 1-hourly weather dynamics for multiple years, we implement the spatiotemporal score function as a massively parallelized convolution operation through time that efficiently manages the available memory—optionally on multiple devices. This extension of the original implementation has proven indispensable for the problem scale. The corresponding Python implementation is provided at https://github.com/schmidtjonathan/Climate2Weather.
@@
The generative model described above can generate random sequences of weather-like patterns. One can think of it as a physical prior in that it represents a model for the general spatial and temporal patterns observed in high-resolution weather dynamics. To tackle the downscaling task specifically, we need the second component (II.) of the SDA method, which allows us to impose the constraints provided by the coarse ESM simulations onto this prior model. This is achieved by conditioning the trained score model. Concretely, while the unconditioned diffusion model generates samples $p(X(0))$, we want to sample from a posterior $p(X(0)\mid Y)$, instead. To that end, we formulate an observation model $p(Y\mid X(0))$ that relates coarse input $Y$ to $X(0)$. We consider a Gaussian observation model
@@
with observation operator $h(X(0))$ that can be a linear or non-linear function [42, 6]. In our case, $h$ selects from the high-resolution state $X$ every six hours and computes a patch-wise averaging in space. Concretely, we thereby encode the assumption that each grid point at each step in the ESM simulation corresponds to the average of a surrounding spatial region of $16\times 16$ grid points that is observed once every six hours. For our experiments, the observation noise $\mathrm{R}$ is selected by running a simple random search on a range of plausible values. To accelerate model selection, we chose $\mathrm{R}$ based on the model’s performance when recovering small-scale information from a two-day window of coarsened reanalysis data. The resulting observation model is used in all presented experiments. Sampling from $p(X(0)\mid Y)$ amounts to replacing the score function in equation 2 by a posterior score function $\nabla_{X(\tau)}\log p_{\tau}(X(\tau)\mid Y)$. For details, we refer the reader to [42], who establish the concept of "Diffusion Posterior Sampling" and to its extension by Rozet and Louppe [6]. To make the conditioning work efficiently at the problem scales considered in our experiments, we use an approximation in the conditioning mechanism, which is described and derived in section A.6.
@@
    …from the test set, we select two days (starting 07 April 2014 at 04:00 AM) and generate artificial coarse observations that represent ESM projections. The spatial resolution is reduced by a factor of 16 and the temporal resolution by a factor of 6. Specifically,
@@
    The two samples are compared to the reanalysis data from the same time period and evaluated. Variations between the samples imply structured uncertainty.
The predicted results from 3. are analyzed regarding their spatial, temporal, and physical consistency with the reanalysis data, their predictive accuracy, and their uncertainty quantification.
Downscaling CMIP6
As a second step, ESM (CMIP 6) simulations will exchange the artificially coarsened reanalysis data for the final experiment. Instead of artificially spatiotemporally subsampled reanalysis data, we condition the score model on two different ensembles from the CMIP 6 data set: the MPI-HR and the HadGEM3 runs. The experimental setup is adopted exactly from the on-model experiments, aside from the ESM conditioning information and the extended considered time horizon of one full year. Concretely, we downscale the ESM simulations for the considered spatial patch from January 1, 06:00 AM until December 31, 06:00 AM (8 73687368\,7368 736hours) for the year 2014, increasing the spatial resolution from $8\times 8$ ($\sim 100$km) to $128\times 128$ ($\sim 6$km) grid points. For the four high-resolution variables, this corresponds to a total of 4×8 736×(128×128)=572 522 496487361281285725224964$\times$ 8\,736$\times$(128$\times$ 128)=572\,522\,4964 × 8 736 × ( 128 × 128 ) = 572 522 496 predicted values given conditioning information that is coarser by a factor of (16×16)×6=1 536161661536(16$\times$ 16$\times$ 6=1\,536( 16 × 16 ) × 6 = 1 536. For this specific experiment, we used four NVIDIA A-100 GPUs, each of which generated two predictions in parallel. Generating one 1-year sample of hourly downscaled predictions for the $128\times 128$-node region takes around two hours on a single NVIDIA A-100 device.

### v4

@@ Discussion
We introduced a probabilistic approach to joint spatial and temporal downscaling of multiple variables from climate to weather scale. The presented framework revolves around a generative diffusion model, which is trained to learn an implicit representation of the dynamical patterns in reanalysis data and serves as a probabilistic emulator for the forward dynamics model. By conditioning the forward model via an observation model on climate-model output, we can sample from a posterior downscaling distribution. Each sample drawn from this posterior adheres to the established statistical relationship between climate output and fine-scale weather and avoids inconsistencies between time steps and variables, which would lead to physically implausible behavior.
Our approach aligns with a broader trend [41, 42, 19, 43, 44, 45] of replacing computationally expensive simulations with statistical models that emulate earth-system dynamics. Rampal et al. [46] give a topical overview and discussion regarding the use of ML methods for statistical downscaling. In particular, generative-modeling techniques, such as normalizing flows [47], GANs [18, 48], and diffusion models [19, 43, 49] have been emerging as a popular model class. Our model extends existing research [for example 45, 50, 19, 43, 51] by enforcing coherence across spatial, temporal, and variable dimensions. In particular, our predictions are sampled from a joint distribution, avoiding the disconnection that typically comes with sampling sequences of temporally independent states for each individual atmospheric variable. Inconsistencies between downscaled time steps and variables make predicted time series as a whole physically implausible and render the predictions unsuitable for downstream applications that require coherent estimates. We present our framework as one possible technique for such coherence-dependent applications, which could include, e.g., driving dynamical climate impact models and studies of compound events [52], which require inter-variable coherence. A recent related method by Srivastava et al. [53] treats the spatial and temporal domain jointly while focusing on a single variable (precipitation). For a chosen, fixed sequence length, their pipeline generates high-resolution estimates from coarse simulations by separating the task into, first, a deterministic statistical downscaling model, followed by a generative model that introduces probabilistic estimates of high-frequency patterns. A related model class are conditional weather generators [8, 46][4, Chapter 13], which require substantial expertise and resources to implement. By training an unconditional generative model, we differ from previous approaches that integrate the conditioning directly into the training process [19, 45, 49]. In contrast to Harder et al. [42], our method enables soft and uncertainty-aware constraints post-training through posterior inference. Hess et al. [43] exploit the iterative-denoising aspect of diffusion models and initialize the generative process at an only partially perturbed target state. This allows conditioning the model on large-scale information, which the model enhances by replacing the remaining noise with local patterns. Notably, this notion of "conditioning" lacks a clear probabilistic interpretation and is thus fundamentally different from how we use the term throughout this work. Further, existing work focuses on mapping between different reanalysis data sets [42, 54], whereas we explicitly developed a model that allows mapping different climate scenarios to finely resolved time series that predict local weather patterns. Our trained model can be readily re-used for mapping from different coarse predictors to the target resolution without re-training, simply by adapting the observation model and input data accordingly. Diffusion bridges have been proposed for unpaired downscaling of fluid dynamics, though their application to climate-model downscaling remains to be demonstrated [44].
The intended scope of our work is to demonstrate the score-based data assimilation framework by Rozet and Louppe [23] as an elegant and flexible technique for probabilistic downscaling and, perspectively, for other inference problems in the context of atmospheric dynamics. More than outperforming existing per-variable and per-time-step downscaling methods, or establishing our downscaling framework as a new, universally favored approach, we aim to provide a framework for downstream applications that
  1. 1.
    require spatio-temporally coherent time series,
  2. 2.
    consider multiple variables jointly,
  3. 3.
    benefit from flexibly encoding additional prior knowledge about the relation between input and output through the post-training conditioning.
The comprehensive feature set offered by our model complicates direct comparisons with existing methods. Nevertheless, we demonstrate that, according to standard statistical distance metrics, our framework maintains competitive performance when evaluated per time step and per variable against benchmark approaches (see Supplementary Table 1). Further, spectral density comparisons between predictions and reanalysis data are presented in Supplementary Section 2, and a concise assessment of predictive skill for wind-power generation is provided in Supplementary Section 4. Supplementary Figure 11 demonstrates that the model uses learned relationships between the jointly processed variables. Looking ahead, we anticipate that the expanded capabilities of the presented framework will broaden its applicability to a wider range of use cases requiring coherent, high-resolution climate information. Additionally, the modular structure of our approach (Figure˜1) enables the straightforward integration of further, potentially complex, domain-specific knowledge via the observation model in future work. For example, established functional relationships between predicted climate variables and external forcings could be incorporated without retraining the generative model. This modularity allows the trained score model to serve as a reusable foundation for subsequent studies. The diffusion model, including trained weights, and the implementations of the methods and experiments are provided at https://github.com/schmidtjonathan/Climate2Weather.
@@
The main limitation of the presented framework is the increased computational cost that comes with coherent predictions. While we have extended the original method by Rozet and Louppe [23] in terms of scalability, enabling it to process substantially longer trajectories, the temporal and spatial extent of the study is still limited by computational demands. Due to the simultaneous processing of the entire state space, memory requirements remain a limiting factor. This study limits itself to a small region (inspired by Langguth et al. [50]) with a diverse range of orography to establish and validate the methodological framework. However, we believe that long-range teleconnections that affect the sub-region under study are, to some extent, reflected in the regional predictions, given that the training data has been drawn from a Europe-wide reanalysis dataset. This argument is visually supported by embedding high-resolution predictions into a larger spatial context of reanalysis data in Supplementary Section 3 (Supplementary Figures 3 to 6). We conclude that the technique is likely most useful for smaller study regions, for which highly accurate predictions of local dynamics are desired.
In summary, our framework enables joint spatial and temporal downscaling of multiple climate variables, producing coherent, high-resolution scenarios from coarse climate simulations. By decoupling the learning of dynamical patterns from the conditioning on new inputs, the model offers a flexible and efficient tool for inference tasks in meteorology and climate science. This approach facilitates the use of long-term climate projections for local impact studies across multiple timescales.
@@ Methods
We analyze four interrelated atmospheric variables, namely the zonal (uas) and meridional (vas) components of near-surface (10 meter) wind speeds, surface (2 meter) air temperature (tas), and sea level pressure (psl) due to their crucial importance for understanding atmospheric dynamics. They evolve on different spatial and temporal scales, enabling insights into the performance of downscaling methods across different scales.
For training (2006-2013) and evaluation (2014) of the model, we use data from the Consortium for Small-Scale Modeling Reanalysis 6 (COSMO-REA6) data set, a high-resolution reanalysis product for the European domain developed by the German Weather Service (Deutscher Wetterdienst; DWD) [55] with a spatial resolution of approximately $6\penalty 10000\$km and hourly temporal resolution. It serves as the ground truth observational dataset in the perfect-predictor evaluation [37] of the presented methodological framework. The COSMO-REA6 data contains errors for some variables and lacks some observations for the years prior to 2006. In this work, only data from 2006 onwards was used.
We apply our downscaling model to historical model runs from two established general circulation models (GCMs) that are part of the sixth phase of the Coupled Model Intercomparison Project (CMIP6) [3], namely the higher-resolution earth system model (ESM) of the Max–Planck Institute (MPI-ESM1.2-HR) [56] and the high-resolution configuration of the third Hadley Centre Global Environment Model in the global coupled configuration 3.1 (HadGEM3-GC3.1-LM) [57]. Both models have an approximately $100\penalty 10000\$km spatial and 6-hourly temporal resolution.
We restrict our analysis to a spatial subregion that is oriented at the benchmark region proposed by Langguth et al. [50]. The region includes parts of Germany, Switzerland, Austria, and the Czech Republic ($6^{\circ}\,\mathrm{E}-16^{\circ}\,\mathrm{E}$, $46^{\circ}\,\mathrm{N}-52^{\circ}\,\mathrm{N}$ (see Supplementary Figure 1) which results in $128\times 128$ grid points per time point for the reanalysis dataset. We use data for the period 2006-2014, the time range in which the reanalysis and the historical GCM runs overlap. The GCM data is spatially re-gridded to a rotated latitude-longitude grid using bilinear interpolation to match the coordinates and grid type of the reanalysis dataset. As a pre-processing step, we apply a quantile-mapping procedure to the GCM data. This mitigates biases by adjusting the value distribution of the GCM to align better with the distribution of the reanalysis product. For this, we use the quantile-mapping implementation provided by the python-cmethods [58] Python package with default parameters.
@@
We approach statistical downscaling as a Bayesian-inference problem [59, 60]: An uncertain estimate for an unknown quantity (here: high-resolution weather) is obtained by first formulating a prior distribution that encodes known, assumed, or learned properties of the unknown quantity. The prior defines the output space of the prediction model. Through an observation model (or "likelihood"), the prior is conditioned on the available information (here: coarsely simulated climate output) to yield the posterior distribution. The posterior ideally captures the uncertainty that is both inherent in the prior and which arises from incomplete information and model mismatch.
For inference in the context of dynamical systems, it is useful to formulate a prior model that somehow represents a mechanistic system that we assume to underlie the unknown dynamics. The core of the presented methodological framework is a score-based diffusion model (DM) [61, 20, 21, 62, 63], which constitutes the prior model of the framework. DMs are an instance of generative models that involve training a deep neural network on a finite set of data points, which can then be used to generate new samples from the underlying data distribution. The basic formulation of DMs involves two main components. Firstly, a forward diffusion process transports the data distribution to a known, tractable distribution, such as a standard normal distribution [20, 21]. In this process, structured data points (like RGB images, weather states, etc.) are successively perturbed until all signals are entirely replaced by noise. The pivotal insight is that a certain class of diffusion processes are known to have a reverse counterpart, providing a generative process that successively refines samples from the noise distribution into structured data. The reversal of the diffusion process requires access to the score function, which can be understood as a function that, for any given degree of perturbation during the reverse diffusion process, separates noise from the signal to remove it. An exemplary such de-noising process for sequences of meridional wind speeds is visualized in Supplementary Figure 10. In the following, we will give a brief overview of the general framework and the extensions to it necessary to obtain the results presented in this work.
@@
The state $X(\tau)$ at $\tau=0$ is a data point $X_{0}\sim\mathcal{D}$ selected from a data set $\mathcal{D}$. The data point is successively perturbed through the diffusion process and thus loses all structure as $\tau\rightarrow\mathrm{T}$. Note that $\tau$ is sometimes referred to as "time", which does not mean physical time but rather a continuous degree of perturbation of the initial state. Steps in physical time will later be denoted in the subscript, e.g., $X_{1:L}(\tau):=(X_{1}(\tau),\dots,X_{L}(\tau))$ will denote a time series of length $L$ perturbed according to $\tau$. The drift $\mathrm{F}$ and dispersion $\mathrm{L}$ define functional properties of the forward process, which is driven by Brownian motion $W(\tau)$. At a final time step $\mathrm{T}$ (often $\mathrm{T}=1$), the process converges to a Gaussian distribution such that $X(\mathrm{T})\sim\mathcal{N}\left(0,\mathrm{\Pi}\right)$, where the final-step covariance $\mathrm{\Pi}$ depends on the choice of $\mathrm{F}$ and $\mathrm{L}$ and is often modeled to be the identity matrix. From a result by Anderson [65], the reverse process of Equation˜2 is known and given as
@@
This process is driven by reverse Brownian motion $\overleftarrow{W}(\tau)$ and depends on the gradient of the log-marginal density, called the score, of the diffused state $X(\tau)$ at every perturbation-time point $\tau$. We call Equation˜3 the generative process as it allows sampling unseen data points from the data distribution simply by
@@
    numerically simulating Equation˜3 backwards through time, starting from $X(\mathrm{T})$ and ending at a generated sample $X(0)$.
There is much existing and active research regarding the sampling algorithm used in step 2., which, in general, is slower and more complex when comparing diffusion models to other methods from the generative-model class [66, 67, 68]. Hess et al. [43] employ a recent extension to the diffusion-model framework that allows single-step sampling. This work uses a standard technique that solves an ordinary differential equation related to the marginal distribution of Equation˜3 [22, 21, 63] following the original proposal by Rozet and Louppe [23].
In practice, the score of $p_{\tau}(X(\tau))$ is not accessible and has to be estimated. It is common practice to train a noise-dependent neural network $s_{\theta}(X(\tau),\tau)$ on a finite set of samples from the data distribution to approximate the true, intractable score function. Most practical approaches optimize a de-noising score-matching objective [69, 20, 22, 21, 62, 63]. In light of this, simulating Equation˜3 is commonly perceived as iteratively de-noising the initial Gaussian random state $X(\mathrm{T})$ and successively refining it to a noise-free data point. For spatial data, a time-conditioned U-Net [70, 20] architecture has proven effective in modelling the score function. The U-Net architecture consists of multiple levels, each containing blocks of convolutional layers and skip connections. Along these levels, the spatial dimensionality of the input data is first successively reduced and then increased again. This encoder-decoder structure results in a bottleneck layer that forces the model to extract a limited set of meaningful features from the input. At each level, skip connections between equal-resolution blocks are introduced to facilitate the learning task by maintaining information throughout the encoding- and decoding process. This work uses a U-Net architecture with a total number of around $72$ million parameters, including a self-attention mechanism at the $8\times 8$-bottleneck layer [20, 27].
The framework used in the presented experiments leverages the work by Rozet and Louppe [23] that extends the basic DM framework, enabling robust sampling from a posterior distribution and generating arbitrary-length sequences of spatiotemporally coherent state trajectories. The idea behind score-based data assimilation (SDA) is to separate an inference task into two parts:
@@
    learning a generative model that generates spatiotemporal patterns from the target output space and
@@
    conditioning the prior model using a functional or statistical constraint (the observation model) that defines the specific inference task.
This separation aligns with the idea of Bayesian inference, in which a posterior estimate is obtained from combining a prior model (I.) with an observation model (II.) through conditioning. A high-level introduction of both parts follows below. For more details regarding the score-based data assimilation method, an extensive evaluation, and the corresponding code base, we refer to the original publication by Rozet and Louppe [23].
@@
The generative model (I.) constitutes the prior in the present Bayesian-inference approach to statistical climate downscaling. As described above, the diffusion model requires a trained score model $s_{\theta}(X(\tau),\tau)$ to iteratively refine an initial random-noise sample into a generated data point (cf. Equation˜3). In our case, the generative model outputs uninformed sequences of high-resolution weather patterns. To generate sequential data, Rozet and Louppe [23] propose to learn a de-noising model on short fixed-length time windows. During training, the time window is flattened into the channel-dimension of the convolutional neural network that represents the score model $s_{\theta}$. This allows the model to learn correlations between different variables at different points in time. In the sampling routine, each point in time is de-noised by computing the score in the context of a surrounding temporal context window of size $w=2k+1$. Rozet and Louppe [23] argue that many dynamical systems are (approximately) Markovian and in order to predict the current state, it is sufficient to regard the Markov blanket of order $k$ around the current time step, instead of the entire trajectory. This yields output dynamics that are temporally coherent without the need for training the network on long sequences, which would not only significantly increase the computational complexity but also fix the desired sequence length $L$ as a pre-determined hyperparameter, greatly limiting flexibility. Through an intricate convolution-like routine, the windows are assembled into trajectories of arbitrary length $L$. The corresponding pseudo-code is taken from Rozet and Louppe [23] and given in Algorithm˜1.
Algorithm 1 Composing the sequential score function $s_{\theta}(X_{1:L}(\tau),\tau)$
1: Input: sample $X_{1:L}(\tau)$ of sequence length $L$, perturbation level $\tau$; Markov order $k$; trained score model $s_{\theta}$.
2: Output: Score $s_{1:L}\approx\nabla_{X_{1:L}(\tau)}\log p_{\tau}(X_{1:L}(\tau))$ for every time step $t\in 1{:}L:=(1,\dots,L)$
3: function $s_{\theta}(X_{1:L}(\tau),\tau)$
4:   $s_{1:k+1}\leftarrow s_{\theta}(X_{1:2k+1}(\tau),\tau)[:\!k+1]$
5:   for $t=k+2$ to Lk1L-k-1 do
6:    $s_{t}\leftarrow s_{\theta}(X_{t-k:t+k}(\tau),\tau)[k+1]$
7:   $s_{L-k:L}\leftarrow s_{\theta}(X_{L-2k:L}(\tau),\tau)[k+1\!:]$
8:   return $s_{1:L}$
@@
For our experiments, we require scaling the original SDA framework to allow for the prediction of very long time horizons with thousands of steps. Unfortunately, sampling a spatiotemporally coherent trajectory in the described manner comes at an increased computational cost and, in particular, significant memory demands since the entire state trajectory must be held in memory. To make the approach scale to predicting 1-hourly weather dynamics for multiple years, we implement the spatiotemporal score function Algorithm˜1 as a massively parallelized convolution operation through time that efficiently manages the available memory—optionally on multiple devices. This extension of the original implementation has proven indispensable for the problem scale. The corresponding Python implementation is provided at https://github.com/schmidtjonathan/Climate2Weather.
@@
The generative model described above can generate random sequences of weather-like patterns. One can think of it as a "physical prior" in that it represents a statistical model for the general spatial and temporal patterns observed in high-resolution weather dynamics. Note that the physical laws are only implicitly represented by this model, through learning it from existing simulations, and no explicit mechanistic laws are encoded in the model. To tackle the downscaling task specifically, we have to inform the prior model about the climate constraints. To this end, we introduce the conditioning mechanism, which allows us to impose the constraints provided by the coarse ESM simulations onto the prior model.
We proceed to describe the conditioning mechanism (II.). While the unconditioned diffusion model generates samples $p(X(0))$, the goal is to sample from a posterior $p(X(0)\mid Y)$, instead. We denote the high-resolution output with $X(0)$ and the coarse input with $Y$. Here, we include the noise level $\tau$ of the sample $X(\tau)$ into the notation (in the parentheses), as it will become relevant in this section. Recall that $X(0)$ is the generated, noise-free sample. An observation model $p(Y\mid X(0))$ relates the coarse input $Y$ to $X(0)$. We consider a Gaussian observation model
@@
with an observation operator $h(X(0))$ that can be a linear or non-linear function [71, 23] and observation- or sensor noise defined by a positive-definite matrix $\mathrm{R}$. In our case, $h$ selects from the high-resolution state $X$ every six hours and computes an area averaging in space (see Equation˜1). For our experiments, the observation noise $\mathrm{R}$ is selected by running a simple random search on a range of plausible values. To accelerate model selection, we chose $\mathrm{R}$ based on the model’s downscaling performance on a two-day window. The resulting observation model is used in all presented experiments. Sampling from $p(X(0)\mid Y)$ amounts to replacing the score function in Equation˜3 by a posterior score function $\nabla_{X(\tau)}\log p_{\tau}(X(\tau)\mid Y)$. The posterior score is obtained through Bayes’ rule by noting that
$\nabla_{X(\tau)}\log p(X(\tau)\mid Y)=\nabla_{X(\tau)}\log p_{\tau}(X(\tau))+\nabla_{X(\tau)}\log p(Y\mid X(\tau)),$ (5)
The normalization constant is not dependent on $X$ and thus vanishes when taking the gradient. The prior score is modeled with the parametric score model $s_{\theta}(X(\tau),\tau)$. Note that the likelihood term $p(Y\mid X(\tau))$ is defined on the perturbed state $X(\tau)$ and has to be approximated to be compatible with the observation model $p(Y\mid X(0))$. For details, we refer the reader to [71, Section 3.1], who establish the concept of "Diffusion Posterior Sampling" and to its extension by Rozet and Louppe [23] (Section 3.2). Further, a mathematical description is given in Supplementary Section 6. To make the conditioning work efficiently at the problem scales considered in our experiments, we introduce an approximation in the conditioning mechanism that avoids the computation of gradients with respect to the score network. This is described and derived in Supplementary Section 6.
@@
    …from the test set, we select an evaluation period and generate artificial coarse observations that represent ESM projections. The spatial resolution is reduced by a factor of $16\times 16$ and the temporal resolution by a factor of 6 (see Equation˜1). Specifically,
@@
    The samples are compared to the reanalysis data from the same time period and evaluated. Variations between the samples provide structured uncertainty.
[…]
Data Availability
The reanalysis data were taken from the COSMO REA-6 reanalysis product [55]. The CMIP6 data [3] was downloaded from the Earth System Grid Federation (ESGF) at https://esgf-node.llnl.gov/projects/cmip6/.